### v1

# Phonon fluctuation diagnostics: Origin of charge order in AV3Sb5 kagome metals

Abstract
The microsopic origin of the charge-density wave (CDW) in AV3Sb5 (A = K, Rb, Cs) kagome metals remains a longstanding question, often revolving around electron-phonon coupling and purely electronic mechanisms involving Van Hove scenarios, nesting, and sublattice interference. To reveal the processes driving the CDW transition, we combine ab-initio calculations analysis of the phonon self-energy and angle-resolved photoemission spectroscopy (ARPES). Our momentum-resolved study, supported by ARPES data, reveals that lattice instabilities in the V-135 family of kagome metals appear to also be driven by electronic states far from high-symmetry points, where these states exhibit the strongest coupling with the phonon modes responsible for the CDW distortion. Footing on an interpretation scheme based on phonon fluctuation diagnostics, our work challenges and revises theories that so far have exclusively attributed CDW formation to nesting effects close to the Fermi level.
Introduction – Kagome metals have been an emerging ground for exotic states of electronic matter [1, 2, 3]. Instabilities in kagome metals are hypothesized to originate in simplified theoretical models from an intricate interplay of topology, electron-electron interactions, and enhanced availability of states due to Van Hove singularity (VHS) at the Fermi level [4, 5, 6, 7, 8, 9, 10]. Yet, kagome metals are real materials featuring higher degree of complexity than what is encoded in few-bands tight-binding fermionic-only models, so that the identification of electronic degrees of freedom relevant for the various instabilities remains a major outstanding point in the field. Here, we identify the electronic degrees of freedom responsible for the lattice instabilities in kagome metals directly. We find that these instabilities, across a series of compounds, involve electronic states deep inside the Brillouin zone (BZ), i.e., away from high symmetry points. This ab-initio based finding challenges current understanding based on low-energy VHS scenarios, nesting, and sublattice interference for Fermi surface instabilities [4, 5, 6, 7, 8]. Capable to adequately tracking down electronic degrees of freedom far of the Fermi level, our findings shed new light on the understanding of recent angle-resolved photoemission spectroscopy (ARPES).
Figure 1: (a) Primitive unit cell of AV3Sb5. (b) CsV3Sb5 band structure along path shown in Fig. 2. The highlighted region indicates the active subspace $\mathcal{A}$ used for the cDFPT calculation. (c) Phonon dispersion of CsV3Sb5 with instability mode at $\bm{q}=\mathrm{M}$ and $\bm{q}=\mathrm{L}$ obtained from DFPT. (d) Phonon dispersion from cDFPT and “unscreening” the DFPT result with the approximate ($g^{\mathcal{A}}g^{*}\rightarrow|g|^{2}$) phonon self-energy.
The family of AV3Sb5 (A = K, Rb, Cs) kagome superconductors has been heavily investigated due to its interesting lattice, frustration, topology, quantum geometry, chiral ordering, time-reversal symmetry breaking, flat bands, and the occurrence of a  charge-density wave (CDW) [11, 12, 13]. In particular, the origin of the charge ordering instability and its connection to superconductivity are currently under intense discussion [14]. The members of the AV3Sb5 group show a CDW transition below 78 K–103 K, where the structure undergoes a $2\times 2\times 2(4)$ reconstruction [15, 16, 17, 14]. At the level of a single layer (2D), the compound is assumed to adopt a Star of David (SD) or an inverse Star of David (ISD) configuration. Experimentally, the 135 V-based kagome exhibit different 3D ground-state reconstructions, where SD and or ISD can be stacked with or without a $\pi$-shift [18, 19]. The rearrangement can be directly linked to the theoretically observed phonon instabilities around the M- and L-points. We will put a particular focus on $\bm{q}=\mathrm{L}$; this is because we find it to exhibit the largest imaginary phonon frequency (Fig. 1(c)), even though either combinations of M and L can in principle contribute to the $2\times 2\times 2$ reconstruction (see Supplementary Information (SI) for $\bm{q}=\mathrm{M}$ analysis).
The structure of V-based 135 kagome compounds is built up by a vanadium kagome lattice (red Fig. 1(a)), surrounded by Sb atoms from top and bottom as well as inside the kagome hexagon. These V3Sb5 layers are separated by the A-atoms and determine the V d and Sb p dominant orbital contribution to the electronic structure around the Fermi level. It is this multi-orbital behavior complicating the understanding of the low-energy physics, and rendering a simple one-orbital kagome description insufficient [4, 20].
Instead of building toy models with specific orbital character to investigate the lattice instabilities, we pinpoint the electronic processes and the related electronic states based on an ab-initio approach derived from density-functional theory (DFT) and density-functional perturbation theory (DFPT). This analysis is performed by examining the $\bm{k}$-momentum-resolved phonon self-energy that can be directly connected to the microscopic origin of the CDW, which we refer to as phonon fluctuation diagnostics [21, 22]. We express the phonon self-energy as a product of the electron susceptibility and electron-phonon coupling, resolved in electronic momentum space. Therefore, phonon fluctuation diagnostics allows to discriminate electronic against electron-phonon contributions to a given lattice instability. In addition, our comparison of fluctuation diagnostics with ARPES measurements provides an unprecedented confirmation of our findings.
The hitherto common narrative with regard to the origin of the CDW mechanism triangulates between the central themes of electronic interactions, phonon modes and electron-phonon coupling, as well as properties of Fermiology such as the nature and Fermi level distance of VHS, Fermi surface nesting, and sublattice interference [4, 5, 6, 7, 8]. One scenario of choice derives the charge order instability from the large density of states of the VHS at the M- and L-points close to the Fermi level. The phonon instability vectors $\bm{q}=\mathrm{M}$ and $\bm{q}=\mathrm{L}$ link these high-symmetry points, suggesting a Peierls mechanism resulting in a $2\times 2$ instability [4]. This explanation suggests a correlation between charge order and VHS proximity to the Fermi level, which is challenged by kagome CsTi3Bi5. There, the energy of the VHS can be tuned to the Fermi level with doping, yet no CDW is observed [23]. Even more drastically, recent findings reveal an anticorrelation between the CDW order parameter and the distance of the VHS to the Fermi level. Setting aside subtleties in the ordering propensities of charge vs nematic order [24] and its substantiation in the distinction between ordinary and higher order VHS, it suggests that the electronic states at the VHS might not be the single driving mechanism for the CDW [25]. Another related scenario centers around more general (Fermi surface) nesting, where, instead of specific high-symmetry points, entire lines in the BZ are connected by the instability vectors [5]. As much as this rational idea could explain the origin of the CDW for an idealistic system with perfect nesting, it struggles to account for real band structures under pressure and strain, where the Fermi surface nesting becomes significantly worse yet the charge order instability persists largely unaffected. While beforementioned scenarios imply a leading role of electronic correlation effects with an at best largely coarse-grained modelling of electron-phonon coupling, the charge order instability could also be determined by a strong or $\bm{k}$-selective electron-phonon coupling [6].
In this article, we seek to shed new light on the origin of the CDW instabilities in AV3Sb5 by taking the modelling of phononic impact on charge order to the next level from an ab-initio perspective. We find the nature of electron-phonon coupling to constitute the key component to the driving mechanism of charge order, and reach suggestive agreement between theoretical modeling and ARPES results.
Instabilities and phonon self-energies – To investigate the origin of the charge density wave (CDW), we first establish the connection between the phonon self-energy $\Pi$ and the lattice instability. The interatomic force constants, and thus the phonons, are largely determined by the response of the electron density to ionic displacements. In the following, we show that the most relevant contributions come from the low-energy electrons with energies close to the chemical potential. We denote this active subspace of low-energy electronic states as $\mathcal{A}$. The contribution of processes within this subspace to the force constants reads
$\Pi_{\bm{q}\mu\nu}^{\mathcal{A}}=\frac{2}{N}\sum_{\bm{k}mn}^{\mathcal{A}}g_{\bm{q}\mu\bm{k}mn}^{\mathcal{A}}\frac{f(\varepsilon_{\bm{k}+\bm{q}m})-f(\varepsilon_{\bm{k}n})}{\varepsilon_{\bm{k}+\bm{q}m}-\varepsilon_{\bm{k}n}}g_{\bm{q}\nu\bm{k}mn}^{*},$ (1)
with the Kohn-Sham energies $\varepsilon$, the corresponding occupations $f$, the (partially) screened electron-phonon coupling $g^{(\mathcal{A})}$, the electronic and phononic wave vectors $\bm{k}$ and $\bm{q}$ and band indices $m,n$ and $\mu,\nu$, and the number of wave vectors summed over $N$. While the summation is constrained to active states, $\{|\bm{k}m\rangle,|\bm{k}{+}\bm{q}m\rangle\}\subset\mathcal{A}$, the partially screened coupling $g^{\mathcal{A}}$ excludes screening from scattering between such states. We utilize an active window $\mathcal{A}$ of 0.1 eV around the Fermi level (indicated in Fig. 1(b)) and compute $\varepsilon$, $g$, and $g^{\mathcal{A}}$ using DFT, DFPT [26], and cDFPT [27].
Figure 2: (a),(b) and (c) The $\bm{k}$-resolved electronic contributions to the phonon self-energy of AV3Sb5 with A = (K, Rb, Cs) for $\bm{q}=\mathrm{L}$, visualized in the $k_{z}=0$ plane. The Fermi contour inside the electronic BZ is shown in red. All members of the vanadium-135 kagome family exhibit similar self-energies with the main contributions occurring in the region between the M-points, however not directly at the M-point. The high-symmetry points and the in-plane component of the utilized $\bm{q}$ are shown in (b).
Compared to the unstable DFPT result in Fig. 1(c) that agrees well with the extensive literature, the cDFPT approach with partially screening inside $\mathcal{A}$ stabilizes the structure (see Fig. 1(d)) [28, 29]. Similarly, by renormalizing the DFPT frequencies with the phonon self-energy through
$\tilde{\omega}_{\bm{q}\mu\nu}^{2}=\omega_{\bm{q}\nu}^{2}\delta_{\mu\nu}+2\sqrt{\omega_{\bm{q}\mu}\omega_{\bm{q}\nu}}\,\Pi_{\bm{q}\mu\nu},$ (2)
one can also avoid the phonon instabilities and reproducing the cDFPT result very accurately [21]. Here, $\tilde{\omega}_{\bm{q}\mu\nu}$ and $\omega_{\bm{q}\nu}$ are the renormalized and ab-initio phonon frequencies.
Figure 3: (a) Fermi surface of KV3Sb5 in the low-temperature phase with the path in red trough adjacent M-points. (b) ARPES band structure below and above the CDW temperature transition overlayed with DFT bands the unfolded DFT band structure of staggered ISD $2\times 2\times 2$ phase and the bands of the primitive cell. Below the CDW transition the bands split up along the identified path between the M-points, as indicated with red arrows.
The stabilization highlights that the CDW instability is encoded in the phonon self-energy or respectively the reduction in screening. The only phonon mode significantly altered by these approaches is the unstable one. This further emphasizes the direct connection of the phonon self-energy with the CDW instability. Due to the agreement in Fig. 1(d), we approximate Eq. 1 with two screened $g$ for simplicity (as $g^{\mathcal{A}}\rightarrow g$ for $\mathcal{A}\rightarrow\emptyset$) [30, 31].
After establishing the critical role of the phonon self-energy regarding the CDW instability in 135-V kagome structures, a comparison between the three members of the family is drawn. Based on ab-initio results, the $\bm{k}$-resolved electronic contributions to the phonon self-energy at the wave vector $\bm{q}=\mathrm{L}$ of the instability is shown in Fig. 2 (Eq. 1 before $\bm{k}$ summation). For clarity, we limit the analysis to one of three equivalent $\bm{q}=\mathrm{L}$ vectors, which causes the features to appear with only C2 symmetry. Nevertheless, with all three $\bm{q}=\mathrm{L}$ the investigated properties retain the full C6 symmetry of the structure. Notably, the $\bm{k}$-dependence of the phonon self-energy is similar across all three materials, indicating a common origin and a consistent behavior of the CDW. The highest amplitudes appear in the regions between two adjacent M-points, which are connected by the in-plane component of the examined $\bm{q}$ vector associated with the instability. However, it is important to emphasize that this range does not extend fully to the M-point, where only minor phonon self-energy contributions are observed. This observation deems the VHS scenario unlikely as the dominant driver of the charge instability. Under this scenario, one would expect strongest phonon self-energy contributions near the M-points, where the density of states diverges.
[…]
The significance of the region between the M-points is further emphasized by comparing unfolded DFT band structures in the energetically favored $2\times 2\times 2$ staggered ISD supercell with ARPES maps. This is done for KV3Sb5 as it was previously reported to exhibit the theoretically expected $2\times 2\times 2$ reconstruction at low temperatures [18, 32]. Both experimentally and theoretically, band splittings can be observed in Fig. 3 along the indicated path through adjacent M-points as the temperature is reduced below the CDW transition temperature. As indicated in Fig. 3(b), these splittings not only occur at the M-point but also between them, consistent with previous reports, further contradicting a dominant VHS scenario [5]. Additionally, the splitting is evident in the significant downshift of the intensity near the Fermi level around M. By combining phonon fluctuation diagnostics with ARPES we experimentally confirm the region in momentum space, which we theoretically identified from the analysis of the phonon self-energies role.
Phonon fluctuation diagnostics –
Figure 4: The $\bm{k}$-momentum-resolved phonon fluctuation diagnostics of the instability mode at $\bm{q}=\mathrm{L}$ for $k_{z}=0$. (a) The electronic susceptibility $\chi^{0}$, (b) the coupling matrix elements $2\omega g^{2}$ (el-ph), and (c) the phonon self-energy $2\omega\varPi$ are shown. The solid red lines indicate the Fermi surface and the dashed lines Fermi surface shifted by $q=\mathrm{L}$ inside the electronic BZ. The strong electron-phonon coupling between the $M$ points dominates the phonon self-energy.
In the following, the phonon self-energy will be further investigated. To disentangle the electronic response and matrix elements, we consider the momentum-resolved phonon self-energy $\Pi_{\bm{q}\mu\nu\bm{k}mn}$ as a product of the electron-phonon coupling $g^{2}_{\bm{q}\mu\nu\bm{k}mn}$ and the electronic susceptibility $\chi^{0}_{\bm{q}\bm{k}mn}$. This allows to distinguish the different contributions of the electronic density response to atomic displacements and enables us to identify the electronic processes behind the phonon softening associated with the origin of the CDW. We representatively discuss KV3Sb5 and refer to the SI for the other compounds. The relevant formula are:
$\Pi_{\bm{q}\mu\nu\bm{k}mn}=g^{2}_{\bm{q}\mu\nu\bm{k}mn}\times\chi^{0}_{\bm{q}\bm{k}mn},$ (3)
$g^{2}_{\bm{q}\mu\nu\bm{k}mn}=g_{\bm{q}\mu\bm{k}mn}^{(\mathcal{A})}g_{\bm{q}\nu\bm{k}mn}^{*}.$ (4)
$\chi^{0}_{\bm{q}\bm{k}mn}=\frac{f(\varepsilon_{\bm{k}+\bm{q}m})-f(\varepsilon_{\bm{k}n})}{\varepsilon_{\bm{k}+\bm{q}m}-\varepsilon_{\bm{k}n}},$ (5)
We analyze the phonon mode at $\bm{q}=\mathrm{L}$ where the highest imaginary frequency can be found, which is indicative of the instability in the system (phonon branch indices $\nu=\mu=\text{soft mode}$). The corresponding k-resolved electron-phonon coupling is obtained with ab-initio calculations and depicted in Fig. 4(a) for $k_{z}=0$ (for computational details, see Methods). We find the strongest electron-phonon coupling in the area between the M-points connected by the in-plane component of the chosen q vector, whereas not at the M-points themselves. To illustrate the electron-phonon coupling coefficient relevant for the phonon self-energy, in Fig. 4(a) we sum over the band indices $m$ and $n$ close to the Fermi level, as only these contribute to the product with the susceptibility (see Eq. 3). The electronic susceptibility evaluated based on Eq. 5 from a Wannier model is shown in Fig. 4(b) together with the $k_{z}=0$ Fermi surface in solid red lines. Additionally, a Fermi surface shifted by the instability wave vector $\bm{q}=\mathrm{L}$ is depicted with dashed lines to visualize possible low-energy transitions. As expected, the strongest contribution to the electron susceptibility can be found around the M-point (see SI for band-resolved susceptibility).
Nevertheless, the electron-phonon coupling only selects the specific k-region between the M-points of the susceptibility for the contribution to the final phonon self-energy in Fig. 4(c). This highlights the dominance of the electron-phonon matrix elements over electronic nesting or Van Hove effects in determining the phonon self-energy. Such a behavior is a strong indicator for electron-phonon-coupling to be the driving factor of the CDW instability.
In the case of a full VHS or general nesting scenario, one would expect the susceptibility to dictate the phonon self-energy as to observe the main contributions around the M-point, where the significant nesting occurs. Such an origin of the instability is in contradiction to our result shown in Fig. 4(c), where the electron-phonon coupling suppresses the susceptibility close to the M-point.
A similar behavior can be observed for various $k_{z}$ cuts and the RbV3Sb5 and CsV3Sb5 compounds (see SI). This suggests a common origin of the CDW for the V-135 kagome group, even though different $2\times 2\times 2$ reconstructions are observed among these compounds, depending on temperature and pressure.
Our findings might crucially contribute to resolving the controversy on the origin of the CDW, were different mechanisms have been proposed. Through disentangling electronic and electron-phonon coupling effects together with a $\bm{k}$-resolved analysis we are able to distinguish between the different scenarios explaining the CDW. Here, our findings suggest electron-phonon coupling as the dominant contribution to the origin of the CDW in AV3Sb5 [33, 34, 35, 36, 37, 6], where electronically driven charge ordering propensity might exhibit a cooperative effect. Our identification of a specific $\bm{k}$-region away from high-symmetry points as a main manifestation regime for the lattice charge order instability is experimentally confirmed by ARPES results. This result leaves room for other many-body effects such as charge-bond order, loop-current order, pair-density waves and superconductivity to originate from the VHS or M-point scattering without dominant phonon contributions. The ultimately complete picture of charge order in kagome metals will likely necessitate aspects of all microscopic sources for charge order formation, which is likely particularly true for the 135 family where electron correlation effects are assumed to be prominent. Our identification of phonon-mediated charge formation away from the BZ location of low energy density of states promises to carry over as a generic theme of all kagome metals. Recently, a crucial importance of phonons was observed in the ScV6Sn6 kagome metal, hinting at the generic significance of phonons in kagome systems [38, 39, 40].
Conclusion – Thanks to the combination of ab-initio calculations and ARPES measurements we largely resolve the long debated question about the origin of the CDW in AV3Sb5 (A = K, Rb, Cs). We utilized the idea of fluctuation diagnostics to investigate the interplay of lattice instabilities with electronic processes and electron-phonon coupling. By directly connecting the phonon self-energy to the lattice instability, we were able to disentangle the Fermiology and electron-phonon matrix elements effects. Here, we identify electron-phonon coupling as the driving effect for CDW within all three V-135 materials.
From our momentum-resolved analysis of the response of the electron density to atomic displacements we find the dominant contributions away from high-symmetry points, and especially the M-point, where the kagome VHS are located. We experimentally confirm that CDW spectral signatures emerge between the M-points away from high-symmetry points. While this does not entirely rule out VHS and ideal nesting scenarios as cause of the predicted and experimentally observed $2\times 2\times 2$ lattice instability, it stresses the unavoidable relevance of phonons in the eventual formation of high-temperature charge order in kagome metals. It will be interesting to transfer our approach to other systems to understand existing or absent lattice instabilities. The kagome compounds with 135 stoichiometry offer an attractive platform for future investigations. This particularly applies to CsTi3Bi5 not exhibiting a CDW [23, 41] and CsCr3Sb5 where different origins of the instability are currently discussed [42].
Acknowledgements – We gratefully acknowledge the Gauss Centre for Supercomputing e.V. (https://www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SuperMUC-NG at Leibniz Supercomputing Centre (https://www.lrz.de). We are grateful for funding support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter ct.qmat (EXC 2147, Project ID 390858490), the University Allowance of the University of Bremen (EXC 2077, Project ID 390741603), FOR 5249 (QUAST, Project ID 449872909 (Project P5), EXC 2056 (Cluster of Excellence “CUI: Advanced Imaging of Matter”, Project No. 390715994) as well as through the Collaborative Research Center SFB 1170 ToCoTronics (Project ID 258499086).